# Taxonomy and Estimation of Multiple Breakpoints in High-Dimensional Factor Models

### v2

@@ 7 Conclusions
This paper introduces a quasi-maximum likelihood (QML) method for estimating multiple breakpoints in high-dimensional factor models. We analyze two distinct types of structural changes and develop corresponding asymptotic theories for the QML estimators. For Type 1 changes, we establish the consistency of the QML estimator, demonstrating that it converges to the true breakpoint with probability approaching one as the sample size diverges. For Type 2 changes, we show that the distance between the QML estimators and the true breakpoints is stochastically bounded.
The proposed QML method is straightforward to implement and computationally efficient, requiring only a single full-sample eigendecomposition. Furthermore, we demonstrate that the number of breaks can be consistently estimated using our novel information criterion. Extensive simulation results confirm the strong finite-sample performance of our approach. Finally, an application to the FRED-MD dataset underscores the method’s practical utility in empirical macroeconomic analysis.

### v3

@@ 7 Conclusions
This paper introduces a quasi-maximum likelihood (QML) method for estimating multiple breakpoints in high-dimensional factor models. We analyze two distinct types of structural changes and develop corresponding asymptotic theories for the QML estimators. For Type 1 changes, we establish the consistency of the QML estimator, demonstrating that it converges to the true breakpoint with probability approaching one as the sample size increases. For Type 2 changes, we show that the distance between the QML estimators and the true breakpoints is stochastically bounded (fractionally consistent).
The proposed QML method is straightforward to implement and computationally efficient, requiring only a single full-sample eigendecomposition. Furthermore, we demonstrate that the number of breaks can be consistently estimated using our information criterion. Extensive simulation results confirm the strong finite-sample performance of our approach. Finally, an application to the FRED-MD dataset underscores the method’s practical utility in empirical macroeconomic analysis.
Supplementary materials
The proofs for the theorems and propositions are given in Appendix A of the supplementary material, the proofs for the technical lemmas are given in Appendix B, sufficient conditions for Assumptions 7 and 8 are provided in Appendix C, additional simulation results are provided in Appendix D, and additional empirical results are reported in Appendix E.
Data availability statement
The data that support the findings of this study are openly available at:
https://www.stlouisfed.org/research/economists/mccracken/fred-databases.